# Scalable Formal Verification via Autoencoder Latent Space Abstraction

### v2

@@ 1. Introduction
Formal verification of continuous-space dynamical systems provides a rigorous way to guarantee that a system satisfies a desired specification (baier2008principles; Tabuada2009; belta2017formal). A standard approach is to express the specification in Linear Temporal Logic (LTL) (baier2008principles) and construct a finite-state abstraction that captures all possible behaviors of the system. Off-the-shelf model-checking tools can then be applied to this abstraction to determine whether the specification is satisfied. This approach, however, faces significant challenges when applied to AI-enabled systems, such as those controlled by neural networks that operate on high-dimensional sensor inputs (e.g., camera images or LiDAR). First, the size of the abstraction grows exponentially with the dimension of the state space. Second, for systems with complex or partially unknown dynamics, computing the transition relation between abstract states becomes highly nontrivial. Yet these challenges are increasingly critical, as AI components are rapidly becoming integral to modern autonomous systems. In this work, we aim to address these challenges by developing a formal verification framework for systems with unknown dynamics that scales to extremely high-dimensional settings.
Model order reduction has been widely used to simplify systems for classical control and verification (obinata2012model; baur2014model; antoulas2001survey). However, for systems with highly complex or unknown dynamics, such reduction techniques often fail to generalize. To address this limitation, modern approaches employ data-driven techniques, such as autoencoders (bengio2013representation), to learn latent representations and latent dynamics for control, typically trading correctness guarantees for flexibility. For systems with unknown dynamics, some works construct finite abstractions with correctness guarantees (jackson2021formal; reed2023promises), but these approaches suffer from scalability limitations. Applying such abstraction techniques directly in a learned latent space is challenging because existing frameworks assume explicit, well-defined system dynamics, whereas latent-space dynamics induced by learned representations may instead be governed by inclusion dynamics, violating these assumptions.
@@
$1.$ Related Works
Latent space control: Many works explore the use of a latent space for planning and control. A common focus is on enabling control for high-dimensional systems such as humanoid robotics or systems that derive control from visual observations. Common techniques for control in the latent space include model predictive control (watter2015embed; masti2021learning), reinforcement learning (hafner2019learning), and Hamilton-Jacobi reachability (nakamura2025generalizing). While these techniques often have great empirical success, they typically lack guarantees on the correctness of the approach which can result in false positives.
Several works have recently appeared to attempt to provide formal guarantees with latent space control, however they often have strong assumptions on the dynamics or can only demonstrate empirical success. In particular, (nakamura2025generalizing; vieira2024morals) make use of autoencoders to define a latent space for control, but lack guarantees on their learned latent space dynamics. (chen2016decomposed; awan2025reduced) propose methods to soundly reason about dynamics in a low-dimensional space, but cannot compose their low-dimensional model with LTL specifications and have strong assumptions on the form of the dynamics. (lutkus2025latent) develops a theory to correctly reason about dynamics in a latent space, but has no method to soundly define the latent domain, preventing the construction of a sound abstraction. We should also stress that various works make use of PCA to identify a low dimensional representation a system for control (matrone2010principal; dalibard2011linear), but as noted in (poupart2002value; bishop2006pattern) without significant structural constraints on the high dimensional representation (e.g. linear dynamics) the latent space identified by PCA is unlikely to preserve the dynamics exactly.
Exceedingly few works attempt to bridge latent space control with formal specifications. (lee2024mission) demonstrates empirical success in driving a system to satisfy an LTL specification though training a latent space controller, but cannot provide correctness guarantees on the mapping to the latent space nor provide guarantees on satisfying the specification post training. (schon2025formal) can construct abstractions in a low dimensional space to verify LTL specifications, but require the original dynamics to be known and decomposable.
@@ 2. Problem Formulation
To specify and reason about temporal properties, we first define a set of atomic proposition $\Pi=\{\mathrm{p}_{1},\ldots,\mathrm{p}_{l}\}$, where $\mathrm{p}_{i}$ is true iff $\mathbf{x}\in r_{i}$. Let $L:X\to 2^{\Pi}$ be a labeling function that assigns to each state the set of atomic propositions that are true at that state. Then, the observation trace of trajectory $\omega_{\mathbf{x}}$ is $\rho=\rho_{0}\rho_{1}\ldots$ where $\rho_{k}=L(\omega_{\mathbf{x}}\!(k))$ for all $k\geq 0$. To express the desired temporal properties, we use Linear Temporal Logic (LTL) (baier2008principles).
Definition 0 (LTL).
@@
The semantics of LTL are defined over infinite traces (baier2008principles) and provided in Appendix A.1. We say infinite trajectory $\omega_{\mathbf{x}}$ satisfies formula $\varphi$, denoted by $\omega_{\mathbf{x}}\models\varphi$, if its observation trace satisfies $\varphi$.
@@
Figure 1. A flowchart of the method used in this paper. From the high-dimensional state space, we gather data. This data is used to train an encoder (Learning Set). We then map the data (Regression Set) and the environment through the encoder, deep kernel GPR is used to learn the latent dynamics from the latent data and the GPR model is composed with the latent environment to create an abstraction. Verification is performed on the abstraction with respect to the specification and results are mapped back to the original environment.:
While Problem 1 may resemble a classical verification problem, the fact that $f$ is both high-dimensional and unknown makes verification highly challenging. To address this challenge, a key element of our approach is to map the high-dimensional system to a lower-dimensional representation using an autoencoder. This reduction, however, must be performed carefully to ensure that the latent (low-dimensional) representation remains a sound abstraction of the original system, thereby preserving the desired formal guarantees. We guarantee a sound abstraction by restricting the class of functions that the autoencoder can represent and reasoning with inclusion dynamics in the latent space, as the mapping in is general not one-to-one.
@@ 3.1. Gaussian Process Regression
Let $\mathcal{D}$ be a given dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}$ obtained from sampling a function of the form $\mathbf{y}=\Psi(\mathbf{x})+\mathbf{v}$, with $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}$. GPR provides a method to predict the output of $\Psi$ at a new point $x^{*}$ using the dataset $\mathcal{D}$. A GP is a collection of random variables, such that any finite collection of those random variables is jointly Gaussian (rasmussen:book:2006). GPR starts by placing a Gaussian prior over $\Psi$ through use of a kernel $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$. Then, under the assumption that $\mathbf{v}$ is a zero-mean Gaussian independent random variable with variance $\sigma_{n}^{2}$ for any new input $x^{*},$ a GP produces a posterior prediction of $\Psi$ conditioned on $\mathcal{D}$ with a mean $\mu_{\mathcal{D}}(x^{*})$ and variance $\sigma^{2}_{\mathcal{D}}(x^{*})$ as:
@@
where $\mathcal{X}=[x_{1},\ldots,x_{m}]^{T}$ and $Y=[y_{1},\ldots,y_{m}]^{T}$ are respectively the vectors of input and output data, $K_{\mathcal{X},\mathcal{X}}$ is a matrix whose $i$-th row and $j$-th column is $\kappa(x_{i},x_{j})$, and $K_{x,\mathcal{X}}=K_{\mathcal{X},x}^{T}$ are vectors whose $i$-th entry is $\kappa(x,x_{i})$. Deep Kernel Learning (DKL) is an extension of GP regression which incorporates a Neural Network $\psi$ into a base kernel as $\kappa(\psi(x),\psi(x^{\prime}))$. DKL that has been shown to enable more accurate mean predictions, smaller posterior variance predictions, and faster abstraction generation (reed2023promises).
@@ 3.1.1. GP Error Bounds
Under the assumption that $\Psi$ is Lipschitz and lives in the Reproducing Kernel Hilbert Space (RKHS) of kernel $\kappa$, then there exists some constant $B>0$ such that $\|\Psi(\cdot)\|_{\kappa}\leq B$ in the compact domain $X$. For many kernels $\kappa$ used in practice, such as the squared exponential kernel, the RKHS is dense in the space of continuous functions (steinwart2001influence).
@@
Proposition 0 ((reed2024error), Theorem 2).
@@ 3.2. Notation
Given a high $n_{x}$-dimensional state $x$, an encoder $h_{enc}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{p}}$ maps $x$ to an $n_{p}$-dimensional state $z=h_{enc}(x)$. We refer to $\mathbb{R}^{n_{p}}$ as the latent space. A decoder $h_{dec}:\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{x}}$ then maps a $z\in\mathbb{R}^{n_{p}}$ to a state $x$. Given a set of points AA, $\text{Conv}(A)$ defines the convex hull of the set. In a metric space $\mathbb{R}^{n_{x}}$, $\mathcal{B}_{r}(x)=\{x^{\prime}\in\mathbb{R}^{n_{x}}\mid\|x-x^{\prime}\|\leq r\}$ denotes a ball of radius $r\geq 0$ centered at $x\in\mathbb{R}^{n_{x}}$.
@@ 4. Finite Abstraction in the Latent Space
We first observe that when the latent space has a lower dimension that the original space, i.e., $n_{p}<n_{x}$, the resulting latent dynamics may be neither injective nor deterministic. In particular, when the dimensionality of System (1) is reduced beyond its minimal realization, i.e., the smallest number of parameters required to exactly reconstruct the observations, the evolution in the latent space is governed by a inclusion dynamics as defined below.
@@ 4.1. Encoder Architecture
Our framework is designed around generating an interpretable latent space where the inclusion dynamics behave well. To this end, we first restrict the encoder to a class of networks that are convex functions, enabling efficient and provably correct reasoning about latent regions of interest (see Section 4.3.1 for details). Convexity is enforced by constraining all weights after the first hidden layer to be nonnegative and by using activation functions that are both convex and monotonic. Networks of this form are known as Input Convex Neural Networks (ICNNs) (amos2017input). Although constraining weights may appear to limit representational power, prior work has shown that ICNNs can be made highly expressive through the use of skip connections. ICNNs are typically employed to enable tractable optimization and inference (sivaprasad2021curious); to the best of our knowledge, their use as an autoencoder has not been explored. Our approach makes use of convexity to map regions to $\mathbb{R}^{n_{p}}$ efficiently.
@@ 4.3. Nondeterministic Transition System Abstraction
A Transition System (TS) is a tuple $\mathcal{N}=(Q,T,L,AP)$, where $Q$ is a finite set of states, $T:Q\times Q\rightarrow\{0,1\}$ is a transition function such that $T(q,q^{\prime})=1$ if $q^{\prime}$ is a successor state of state $q\in Q$, APAP is a set of atomic propositions, and $L:Q\rightarrow 2^{AP}$ is a labeling function that assigns to each $q\in Q$ a subset of APAP. $\mathcal{N}$ is called Non-deterministic (NTS) if $\sum_{q^{\prime}\in Q}T(q,q^{\prime})>1$.
@@
To construct a finite NTS of System (1) in the latent space, we must first define the latent domain and its partitioning to obtain the set of discrete states $Q$. Next, we need a method to label regions of interest in the latent space that correspond to the regions of $R$ to define $L$ and APAP, and a method to identify transitions between discrete states to construct $T$. We begin by discussing how regions in $\mathbb{R}^{n_{x}}$ are mapped into the latent space.
@@ 4.3.1. Mapping Regions of Interest
To reason about the temporal properties of the system in the latent space, it is necessary to map the regions of interest $r\in R$ from $\mathbb{R}^{n_{x}}$ to $\mathbb{R}^{n_{p}}$. To retain guarantees on the correctness of the abstraction, it is necessary to have guaranteed under- and over-approximations of these regions in the latent space. Translating regions down to the latent space can be a challenge when using a generic encoder architecture, and is akin to the neural network verification problem where given an input set we would like to under- or over-approximate the output set of the network, e.g., (wang2021beta). These methods are generally known to work best when given small input sets. Hence to get tight approximations, these methods require partitioning the high-dimensional domain, which could be computationally expensive due to the exponential growth of the size of a partition with the dimensionality. Arbitrary encoder architectures can also induce folding in the domain, resulting in disjoint regions mapping to the same area in the latent space which can make correct labeling a challenge, hence our restriction of the architecture to CICO networks (see Section 4.1).
@@
Given convex set $r_{i}\subset\mathbb{R}^{n_{x}}$, the encoding of this set $r_{Z,i}=\text{Post}_{h_{enc}}(r_{i})$ is also convex. Directly calculating $r_{Z,i}$ is generally infeasible; however, we can sample a set of NN points from $r_{i}$ and map each point through $h_{enc}$ to define a set in $\mathbb{R}^{n_{p}}$ where the convex hull is a subset of $r_{Z,i}$.
We can also over-approximate $r_{Z,i}$ with high confidence through sampling by adapting the method proposed in (lew2022simple), called $\epsilon$-RandUp.
Proposition 0 ($\epsilon$-RandUp (lew2022simple), Theorem 2).
@@
We can formally bound $\epsilon(N,\delta)$ with $L_{h}\left(\frac{v_{n_{x}}}{Nv_{n_{p}}}\log{\frac{(4L_{h}\sqrt{n_{x}})^{n_{x}}}{\delta}}\right)^{1/{n_{p}}}$ where $v_{i}$ is the volume of an $i$-dimensional ball, as under the assumption that $r$ is defined as a subset of a unit ball and is uniformly sampled we can take $D(\partial r,\epsilon/2L_{h})\leq(4L_{h}\sqrt{n_{x}})^{n_{x}}$ and $\Lambda_{L_{h}}^{\epsilon}>\frac{v_{n_{p}}}{v_{n_{x}}L_{h}^{n_{p}}}\epsilon^{n_{p}}$ (lew2022simple). Note that $\epsilon$-RandUp does not require $h_{enc}$ to be a convex function or $r\in R$ to be a convex set, but the convexity of $h_{enc}$ enables us to identify both under- and over-approximations that are arbitrarily tight by increasing the number NN of samples used ((lew2022simple), Theorem 1). A visualization of the process is shown in Figure 2.
We highlight that, through this sampling-based method, the under-approximation of $r_{Z}$ is guaranteed deterministically, whereas the over-approximation holds with high confidence. Alternate methods, e.g., (wang2021beta), can be used for a deterministic over-approximation, though at the cost of reduced efficiency.
Figure 2. Visualization of embedding a region $X_{r}$ into the latent space, with the true boundary of the set shown in black, sampled points $x_{i}$ in cyan, and the Convex Hull under/over-approximations shown in green. :

### v3

@@ 1. Introduction
Formal verification of continuous-space dynamical systems provides a rigorous way to guarantee that a system satisfies a desired specification (Baier and Katoen, 2008; Tabuada, 2009; Belta et al., 2017). A standard approach is to express the specification in Linear Temporal Logic (LTL) (Baier and Katoen, 2008) and construct a finite-state abstraction that captures all possible behaviors of the system. Off-the-shelf model-checking tools can then be applied to this abstraction to determine whether the specification is satisfied. This approach, however, faces significant challenges when applied to AI-enabled systems, such as those controlled by neural networks that operate on high-dimensional sensor inputs (e.g., camera images or LiDAR). First, the size of the abstraction grows exponentially with the dimension of the state space. Second, for systems with complex or partially unknown dynamics, computing the transition relation between abstract states becomes highly nontrivial. Yet these challenges are increasingly critical, as AI components are rapidly becoming integral to modern autonomous systems. In this work, we aim to address these challenges by developing a formal verification framework for systems with unknown dynamics that scales to extremely high-dimensional settings.
Model order reduction has been widely used to simplify systems for classical control and verification (Obinata and Anderson, 2012; Baur et al., 2014; Antoulas et al., 2001). However, for systems with highly complex or unknown dynamics, such reduction techniques often fail to generalize. To address this limitation, modern approaches employ data-driven techniques, such as autoencoders (Bengio et al., 2013), to learn latent representations and latent dynamics for control, typically trading correctness guarantees for flexibility. For systems with unknown dynamics, some works construct finite abstractions with correctness guarantees (Skovbekk et al., 2025; Reed et al., 2023), but these approaches suffer from scalability limitations. Applying such abstraction techniques directly in a learned latent space is challenging because existing frameworks assume explicit, well-defined system dynamics, whereas latent-space dynamics induced by learned representations may instead be governed by inclusion dynamics, violating these assumptions.
@@
1.1. Related Works: Latent space control
Many works explore the use of a latent space for planning and control. A common focus is on enabling control for high-dimensional systems such as humanoid robotics or systems that derive control from visual observations. Common techniques for control in the latent space include model predictive control (Watter et al., 2015; Masti and Bemporad, 2021), reinforcement learning (Hafner et al., 2019), and Hamilton-Jacobi reachability (Nakamura et al., 2025). While these techniques often have great empirical success, they typically lack guarantees on the correctness of the approach which can result in false positives.
Several works have recently appeared to attempt to provide formal guarantees with latent space control, however they often have strong assumptions on the dynamics or can only demonstrate empirical success. In particular, (Nakamura et al., 2025; Vieira et al., 2024) make use of autoencoders to define a latent space for control, but lack guarantees on their learned latent space dynamics. (Chen and Sankaranarayanan, 2016; Awan and Zamani, 2025) propose methods to soundly reason about dynamics in a low-dimensional space, but cannot compose their low-dimensional model with LTL specifications and have strong assumptions on the form of the dynamics. (Lutkus et al., 2025) develops a theory to correctly reason about dynamics in a latent space, but has no method to soundly define the latent domain, preventing the construction of a sound abstraction. We should also stress that various works make use of PCA to identify a low dimensional representation a system for control (Matrone et al., 2010; Dalibard and Laumond, 2011), but as noted in (Poupart and Boutilier, 2002; Bishop and Nasrabadi, 2006) without significant structural constraints on the high dimensional representation (e.g. linear dynamics) the latent space identified by PCA is unlikely to preserve the dynamics exactly.
Exceedingly few works attempt to bridge latent space control with formal specifications. (Lee et al., 2024) demonstrates empirical success in driving a system to satisfy an LTL specification though training a latent space controller, but cannot provide correctness guarantees on the mapping to the latent space nor provide guarantees on satisfying the specification post training. (Schön et al., 2025) can construct abstractions in a low dimensional space to verify LTL specifications, but require the original dynamics to be known and decomposable.
@@ 2. Problem Formulation
To specify and reason about temporal properties, we first define a set of atomic proposition $\Pi=\{\mathrm{p}_{1},\ldots,\mathrm{p}_{l}\}$, where $\mathrm{p}_{i}$ is true iff $\mathbf{x}\in r_{i}$. Let $L:X\to 2^{\Pi}$ be a labeling function that assigns to each state the set of atomic propositions that are true at that state. Then, the observation trace of trajectory $\omega_{\mathbf{x}}$ is $\rho=\rho_{0}\rho_{1}\ldots$ where $\rho_{k}=L(\omega_{\mathbf{x}}\!(k))$ for all $k\geq 0$. To express the desired temporal properties, we use Linear Temporal Logic (LTL) (Baier and Katoen, 2008).
Definition 0 (LTL Syntax).
@@
The semantics of LTL are defined over infinite traces (Baier and Katoen, 2008) and provided in Appendix A.1. We say infinite trajectory $\omega_{\mathbf{x}}$ satisfies formula $\varphi$, denoted by $\omega_{\mathbf{x}}\models\varphi$, if its observation trace satisfies $\varphi$.
@@
Figure 1. A flowchart of the method used in this paper. From the high-dimensional state space, we gather data. This data is used to train an encoder (Learning Set). We then map the data (Regression Set) and the environment through the encoder, deep kernel GPR is used to learn the latent dynamics from the latent data and the GPR model is composed with the latent environment to create an abstraction. Verification is performed on the abstraction with respect to the specification and results are mapped back to the original environment.
While Problem 1 may resemble a classical verification problem, the fact that $f$ is both high-dimensional and unknown makes verification highly challenging. To address this challenge, a key element of our approach is to map the high-dimensional system to a lower-dimensional representation using an autoencoder. This reduction, however, must be performed carefully to ensure that the latent (low-dimensional) representation remains a sound abstraction of the original system, thereby preserving the desired formal guarantees. We guarantee a sound abstraction by restricting the class of functions that the autoencoder can represent and reasoning with inclusion dynamics in the latent space, as the mapping is in general not one-to-one.
@@ 3.1. Gaussian Process Regression
Let $\mathcal{D}$ be a given dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}$ obtained from sampling a function of the form $\mathbf{y}=\Psi(\mathbf{x})+\mathbf{v}$, with $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}$. GPR provides a method to predict the output of $\Psi$ at a new point $x^{*}$ using the dataset $\mathcal{D}$. A GP is a collection of random variables, such that any finite collection of those random variables is jointly Gaussian (Rasmussen et al., 2006). GPR starts by placing a Gaussian prior over $\Psi$ through use of a kernel $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$. Then, under the assumption that $\mathbf{v}$ is a zero-mean Gaussian independent random variable with variance $\sigma_{n}^{2}$ for any new input $x^{*},$ a GP produces a posterior prediction of $\Psi$ conditioned on $\mathcal{D}$ with a mean $\mu_{\mathcal{D}}(x^{*})$ and variance $\sigma^{2}_{\mathcal{D}}(x^{*})$ as:
@@
where $\mathcal{X}=[x_{1},\ldots,x_{m}]^{T}$ and $Y=[y_{1},\ldots,y_{m}]^{T}$ are respectively the vectors of input and output data, $K_{\mathcal{X},\mathcal{X}}$ is a matrix whose $i$-th row and $j$-th column is $\kappa(x_{i},x_{j})$, and $K_{x,\mathcal{X}}=K_{\mathcal{X},x}^{T}$ are vectors whose $i$-th entry is $\kappa(x,x_{i})$. Deep Kernel Learning (DKL) is an extension of GP regression which incorporates a Neural Network $\psi$ into a base kernel as $\kappa(\psi(x),\psi(x^{\prime}))$. DKL that has been shown to enable more accurate mean predictions, smaller posterior variance predictions, and faster abstraction generation (Reed et al., 2023).
@@ 3.1.1. GP Error Bounds
Under the assumption that $\Psi$ is Lipschitz and lives in the Reproducing Kernel Hilbert Space (RKHS) of kernel $\kappa$, then there exists some constant $B>0$ such that $\|\Psi(\cdot)\|_{\kappa}\leq B$ in the compact domain $X$. For many kernels $\kappa$ used in practice, such as the squared exponential kernel, the RKHS is dense in the space of continuous functions (Steinwart, 2001).
@@
Proposition 0 ((Reed et al., 2025), Theorem 2).
@@ 3.2. Notation
Given a high $n_{x}$-dimensional state $x$, an encoder $h_{enc}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{p}}$ maps $x$ to an $n_{p}$-dimensional state $z=h_{enc}(x)$. We refer to $\mathbb{R}^{n_{p}}$ as the latent space. A decoder $h_{dec}:\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{x}}$ then maps a $z\in\mathbb{R}^{n_{p}}$ to a state $x$. Given a set of points $A$, $\text{Conv}(A)$ defines the convex hull of the set. In a metric space $\mathbb{R}^{n_{x}}$, $\mathcal{B}_{r}(x)=\{x^{\prime}\in\mathbb{R}^{n_{x}}\mid\|x-x^{\prime}\|\leq r\}$ denotes a ball of radius $r\geq 0$ centered at $x\in\mathbb{R}^{n_{x}}$.
@@ 4. Finite Abstraction in the Latent Space
We first observe that when the latent space has a lower dimension than the original space, i.e., $n_{p}<n_{x}$, the resulting latent dynamics may be neither injective nor deterministic. In particular, when the dimensionality of System (1) is reduced beyond its minimal realization, i.e., the smallest number of parameters required to exactly reconstruct the observations, the evolution in the latent space is governed by inclusion dynamics as defined below.
@@ 4.1. Encoder Architecture
Our framework is designed around generating an interpretable latent space where the inclusion dynamics behave well. To this end, we first restrict the encoder to a class of networks that are convex functions, enabling efficient and provably correct reasoning about latent regions of interest (see Section 4.3.1 for details). Convexity is enforced by constraining all weights after the first hidden layer to be nonnegative and by using activation functions that are both convex and monotonic. Networks of this form are known as Input Convex Neural Networks (ICNNs) (Amos et al., 2017). Although constraining weights may appear to limit representational power, prior work has shown that ICNNs can be made highly expressive through the use of skip connections. ICNNs are typically employed to enable tractable optimization and inference (Sivaprasad et al., 2021); to the best of our knowledge, their use as an autoencoder has not been explored. Our approach makes use of convexity to map regions to $\mathbb{R}^{n_{p}}$ efficiently.
@@ 4.3. Nondeterministic Transition System Abstraction
A Transition System (TS) is a tuple $\mathcal{N}=(Q,T,L,AP)$, where $Q$ is a finite set of states, $T:Q\times Q\rightarrow\{0,1\}$ is a transition function such that $T(q,q^{\prime})=1$ if $q^{\prime}$ is a successor state of state $q\in Q$, $AP$ is a set of atomic propositions, and $L:Q\rightarrow 2^{AP}$ is a labeling function that assigns to each $q\in Q$ a subset of $AP$. $\mathcal{N}$ is called Non-deterministic (NTS) if $\sum_{q^{\prime}\in Q}T(q,q^{\prime})>1$.
@@
To construct a finite NTS of System (1) in the latent space, we must first define the latent domain and its partitioning to obtain the set of discrete states $Q$. Next, we need a method to label regions of interest in the latent space that correspond to the regions of $R$ to define $L$ and $AP$, and a method to identify transitions between discrete states to construct $T$. We begin by discussing how regions in $\mathbb{R}^{n_{x}}$ are mapped into the latent space.
@@ 4.3.1. Mapping Regions of Interest
To reason about the temporal properties of the system in the latent space, it is necessary to map the regions of interest $r\in R$ from $\mathbb{R}^{n_{x}}$ to $\mathbb{R}^{n_{p}}$. To retain guarantees on the correctness of the abstraction, it is necessary to have guaranteed under- and over-approximations of these regions in the latent space. Translating regions down to the latent space can be a challenge when using a generic encoder architecture, and is akin to the neural network verification problem where given an input set we would like to under- or over-approximate the output set of the network, e.g., (Wang et al., 2021). These methods are generally known to work best when given small input sets. Hence to get tight approximations, these methods require partitioning the high-dimensional domain, which could be computationally expensive due to the exponential growth of the size of a partition with the dimensionality. Arbitrary encoder architectures can also induce folding in the domain, resulting in disjoint regions mapping to the same area in the latent space which can make correct labeling a challenge, hence our restriction of the architecture to CICO networks (see Section 4.1).
@@
Given convex set $r_{i}\subset\mathbb{R}^{n_{x}}$, the encoding of this set $r_{Z,i}=\text{Post}_{h_{enc}}(r_{i})$ is also convex. Directly calculating $r_{Z,i}$ is generally infeasible; however, we can sample a set of $N$ points from $r_{i}$ and map each point through $h_{enc}$ to define a set in $\mathbb{R}^{n_{p}}$ where the convex hull is a subset of $r_{Z,i}$.
We can also over-approximate $r_{Z,i}$ with high confidence through sampling by adapting the method proposed in (Lew et al., 2022), called $\epsilon$-RandUp.
Proposition 0 ($\epsilon$-RandUp (Lew et al., 2022), Theorem 2).
@@
We can formally bound $\epsilon(N,\delta)$ with $L_{h}\left(\frac{v_{n_{x}}}{Nv_{n_{p}}}\log{\frac{(4L_{h}\sqrt{n_{x}})^{n_{x}}}{\delta}}\right)^{1/{n_{p}}}$ where $v_{i}$ is the volume of an $i$-dimensional ball, as under the assumption that $r$ is defined as a subset of a unit ball and is uniformly sampled we can take $D(\partial r,\epsilon/2L_{h})\leq(4L_{h}\sqrt{n_{x}})^{n_{x}}$ and $\Lambda_{L_{h}}^{\epsilon}>\frac{v_{n_{p}}}{v_{n_{x}}L_{h}^{n_{p}}}\epsilon^{n_{p}}$ (Lew et al., 2022). Note that $\epsilon$-RandUp does not require $h_{enc}$ to be a convex function or $r\in R$ to be a convex set, but the convexity of $h_{enc}$ enables us to identify both under- and over-approximations that are arbitrarily tight by increasing the number $N$ of samples used ((Lew et al., 2022), Theorem 1). A visualization of the process is shown in Figure 2.
We highlight that, through this sampling-based method, the under-approximation of $r_{Z}$ is guaranteed deterministically, whereas the over-approximation holds with high confidence. Alternate methods, e.g., (Wang et al., 2021), can be used for a deterministic over-approximation, though at the cost of reduced efficiency.
Figure 2. Visualization of embedding a region $X_{r}$ into the latent space, with the true boundary of the set shown in black, sampled points $x_{i}$ in cyan, and the Convex Hull under/over-approximations shown in green.